OpenStudy (anonymous):

Helium ( ) is adiabaticly expanded (very slowly) in a well insulated container to a volume four times its original volume. It started at room temerature (22 degrees Celsius) and at a pressure of one atmosphere. a. What is the final pressure? b. What is the final temperature? Consider Helium to act as an ideal gas.

OpenStudy (anonymous):

yay thermodynamics

OpenStudy (anonymous):

can you give me original volume :D

OpenStudy (anonymous):

showwy :P helium = 1.67. that is the missing number

OpenStudy (anonymous):

Am I getting warm ? or gone cold? Thermo has never been my thing... sorry dont have a whiteboard..worked out on my work bench...

OpenStudy (anonymous):

for a monoatomic ideal gas undergoing an adiabatic process, PV^(gamma) is constant and gamma is 5/3. But PV = nRT, so P = nRT/V and TV^(gamma-1) [or TV^2/3 for He] is constant. If V goes up by a factor of 4, T must go down by a factor of 4^2/3= 2sqrt2. Given that, just use PV/T = constant to solve for the new P. Remember to use kelvins for T and not celsius....